g ie udta is srted. ou may want to check it for your own knowledge. 4. Make a histogram for these gasoline prices 2:40 to 2:49 2-50 t0 259 260 t269 27010279 280 10 269 290 1o 2.9 2.60 to 2.69 2.70 to 2.79 2.80 to 2.89 | 2.90 to 2.99 5. Subtract the minimum from the data below to have smaller easier numbers to work with: 2.67 2.59 2.59 2.88 2.65 2.44 Le raldesst o 21 lo b.us |D.?|poslal 510-5| 6. Find the variance of the 2.87, 2.99, 2.65, 2.44, 2.87, and 2.85. Assume it is a sample:L
Answer #1

5. Subtract the minimum from the data below to have smaller easier numbers to work with

Answer: 2.87 2.99 2.65 2.44 2.87 2.85 0.41 2.67 2.59 2.59 2.88 2 2 2 2 2 2 2 2 2 2 2 0.43 0.55 0.21 0 0.43 0.23 0.15 0.15 0.446. Find the variance of the 2.87,2.99, 2.65. 2.44,2.87 and 2.85. Asuume it is a sample.

Answer: To find the variance, we have to use the below formula.

Variance = Σ rl

Where Σ x 2.87-2.99 + 2.65+ 2.44-2.87-2.85 6 = 2.78

Therefore the variance is

(2.87- 2.78)(2.99 - 2.78)2(2.65 2.78) (2.44-2.78)(2.87-2.78) (2.85 2.78)2 V ariance -=\frac{0.0081+0.0441+0.0169+0.1156+0.0081+0.0049}{5}=\frac{0.1977}{5}=0.04

Therefore the variance=0.04 rounded to 3 decimals